### v2

# Agent-as-Judge for Factual Summarization of Long Narratives

Abstract
Large Language Models (LLMs) have demonstrated near-human performance in summarization tasks based on traditional metrics such as ROUGE and BERTScore. However, these metrics do not adequately capture critical aspects of summarization quality, such as factual accuracy, particularly for long narratives (>100K tokens). Recent advances, such as LLM-as-a-Judge, address the limitations of metrics based on lexical similarity but still exhibit factual inconsistencies, especially in understanding character relationships and states. In this work, we introduce NarrativeFactScore (NFS), the first “Agent-as-a-Judge” framework that evaluates and refines factuality in narrative summarization. By leveraging a Character Knowledge Graph (CKG) extracted from input narrative, NarrativeFactScore evaluates the factuality and provides actionable guidance for refinement, such as identifying missing or erroneous facts. Our experimental results demonstrate that constructing the CKG enables reasoning with 1/3 of the factuality computation used in prior approach, and achieve three times higher correlation with human judgments. Furthermore, refinement with actionable guidance improves the quality of the summary. [^1] https://github.com/YeonseokJeong/NarrativeFactScore
Agent-as-Judge for Factual Summarization of Long Narratives
Yeonseok Jeong1, Minsoo Kim1, Seung-won Hwang2thanks: Corresponding Author, Byung-Hak Kim3 IPAI, Seoul National University1, Seoul National University2, CJ Group3 {jys3136, minsoo9574, seungwonh}@snu.ac.kr bhak.kim@cj.net
1 Introduction
Figure 1: Comparison of factuality evaluation by LLM and Agent Judge with NarrativeFactScore. Given scenes from The Lord of the Rings, the summary incorrectly claims “Sauron is pursuing Gandalf.” The LLM Judge assigns 100% factuality score, while our Agent Judge correctly identifies this error through analyzing atomic facts about characters, assigning 75% NarrativeFactScore, with specific feedback.
The rise of LLMs OpenAI (2023); Dubey et al. (2024) has brought significant advancements to summarization tasks, achieving performance close to human levels Pu et al. (2023). Most evaluation metrics Lin (2004); Zhang et al. (2019); Yuan et al. (2021) for summarization measure lexical or semantic similarity between summary and ground truth.
In our target scenario of summarizing long narratives (> 100K tokens), metrics such as BooookScore Chang et al. (2024) can measure coherence, but evaluating factuality has remained challenging Subbiah et al. (2024). This is because it requires comparing summaries not only against complex facts but also against the evolving relationships among characters in long narratives. Thus, judging the factuality of such long narratives has therefore inevitably relied on costly human evaluations Kim et al. (2024).
More recently, LLM-as-a-Judge metrics Min et al. (2023); Bishop et al. (2024) have leveraged LLM to assess the factuality, offering a more cost-effective alternative to human annotations. If applied to narrative summarization, these metrics split the summary into smaller units, retrieve similar scenes from the input story, and quantify factuality by LLM.
However, directly using LLM to evaluate factuality has two limitations. First, as demonstrated by Kim et al. (2024), the LLM judge often fails to accurately assess factuality in narratives that require indirect reasoning, such as understanding character relationships or states. For example, in Figure 1, although Sauron is pursuing Frodo in order to obtain the Ring in The Lord of the Rings, the LLM judge inaccurately evaluates the factuality of a summary which incorrectly reports that “Sauron is pursuing Gandalf”. This limitation stems from the inability of the LLM judge to consistently track and reason about character relationships, highlighting the need to maintain structured and consistent CKG.
Second, the LLM judge evaluates factuality as a score with no interpretability, whereas ROUGE Lin (2004) and BERTScore Zhang et al. (2019) enable to interpret why a given score was assigned and further facilitates the generation of actionable feedback. Desirably, evaluation metrics for summary can provide feedback, when the score is low, to explain why it is incorrect and suggest how to improve.
We propose an Agent-as-a-Judge Zhuge et al. (2024) framework, using interpretable evaluation of summaries with a novel NarrativeFactScore, based on which we can refine and improve summary quality. CKG achieves consistency by constructing a names graph that consolidates character aliases and variations across scenes and by performing multiple rounds of relationship extraction, selecting relationships that frequently appear across scenes as edges, inspired by Wang et al. (2023). This construction process ensures that only well-supported character relationships are retained. By leveraging this consistent relationship graph when evaluating the factuality, we can accurately assess even complex narrative facts that require understanding intricate character dynamics.
To improve the interpretability of the metric, NarrativeFactScore also provides feedback for interpretation and refinement when the summary is incorrect. For each statement in the summary, our metric retrieves relevant scenes and character relationships from our CKG to calculate a factuality score. Based on the retrieved evidence, ours evaluates each statement and generates feedback identifying discrepancies between claims and supporting evidence. Since our metric operates autonomously, it is more cost-effective and faster than Human-as-a-Judge. In addition, it offers feedback for low scores, which makes it more reliable than LLM-as-a-Judge metrics. Recognizing the causes of low scores also contributes to generating more accurate summaries through agent-based refinement.
Using NarrativeFactScore provides two key advantages for long narrative summarization. First, it offers a labor-efficient and fast metric that also approximates human evaluation when evaluating the factuality of summaries. Our metric demonstrates a statistically strong correlation with human evaluation, and a test for differences between human evaluation and our metric yielded statistically significant results, with the p-value falling below 0.05. Second, since it provides feedback on factually incorrect parts, agent-based refinement can improve summarization performance. We show that agent-based refinement improves factuality (+14.03), ROUGE (+2.05), and BERTScore (+0.13) on MovieSum Saxena and Keller (2024a), a movie script summarization dataset, and also improves factuality (+12.26), ROUGE (+2.47), and BERTScore (+0.21) on MENSA Saxena and Keller (2024b), a movie scene saliency dataset.
2 Related Work
2.1 Long Narrative Summarization
Summarizing long narratives Saxena and Keller (2024a, b) is challenging due to the high computational and memory demands required by transformer-based models. In prior work Pilault et al. (2020); Li et al. (2021); Wu et al. (2021); Chang et al. (2024), a method called hierarchical merging was introduced, where individual chunks of the narrative are summarized separately and then combined to form a coherent final summary. Although this method preserves the logical structure of the narrative, hallucinations remain a frequent challenge, especially when capturing global information such as character relationships. Thus, our focus is on improving the factuality of the summaries.
2.2 Character Knowledge Graph (CKG)
Since characters are integral to narrative Gurung and Lapata (2024), prior work has aimed to construct a graph to easily utilize them. In narrative texts, CKG shows the unidirectional relationship between a subject and an object character. This process is similar to creating a triple (subject-predicate-object) list in knowledge graph construction Chen et al. (2020). Andrus et al. (2022) utilized the OpenIE system Angeli et al. (2015) for story completion and question-answering tasks, integrating it with GPT-3 Brown et al. (2020) to enhance its effectiveness. Alternatively, a recent method Zhao et al. (2024) that assembles CKG directly using LLMs is more robust approach, as it better captures the nuanced and complex relationships. Our distinction lies not only in constructing CKGs but also in utilizing them to measure and enhance factuality.
2.3 Summarization Metrics for Evaluating Factuality
In recent research, efforts have been made to evaluate factuality of long documents. LongDocFACTScore Bishop et al. (2024) improves this process by calculating BARTScore Yuan et al. (2021) only on the semantically similar portions of the source text for each summary sentence, making it an effective method for handling long documents. FActScore Min et al. (2023) further enhances factuality evaluation by decomposing text into atomic facts and verifying each with LLM using information retrieved from the knowledge source. Unlike these metrics, our metric focuses on character relationships to accurately evaluate factuality and provide actionable feedback to refine factually incorrect parts.
3 Proposed Method
Figure 2: The main figure illustrates the overall process of evaluation and refinement, which includes three main stages. First, it shows the extraction of CKG $G$ from narrative $\mathcal{N}$. Next, it depicts the calculation of factuality by comparing the decomposed summary $a_{k}$ against the retrieved character relationship subgraph $g$ and narrative scene $\mathcal{S}_{i}$. Finally, it illustrates the agent-based refinement process, where feedbacks ($f_{1},f_{2},$...) are used to improve the factual accuracy of the summary.
In this section, we elaborate on NarrativeFactScore for evaluating factuality of long narrative summarization. Figure 2 illustrates three phases of our framework, which will be detailed in Section 3.1, 3.2, and 3.3 respectively.
3.1 CKG Extraction
We construct a consistent CKG, to overcome the inconsistencies of CKG reported in Kim et al. (2024); Zhao et al. (2024), losing information Liu et al. (2024) in long narratives and failing to reason over many implicit relationships at once. To address these issues, we perform reasoning multiple times Wang et al. (2023) for each scene and select frequent relationships to improve consistency and accuracy. We note this requires only 1/3 of the original cost while improving correlation threefold. (See Appendix 5.4.)
Given a narrative represented as a collection of scenes $\mathcal{N}=\{\mathcal{S}_{1},\mathcal{S}_{2},\ldots,\mathcal{S}_{m}\}$, where $m$ denotes the number of scenes, the goal is to extract a graph $G$ that encapsulates character relationships. Each scene $\mathcal{S}_{i}\ (1\leq i\leq m)$ is processed individually to extract relation triples (subject-predicate-object) using GPT-4o-mini OpenAI (2023), as detailed in Section D.1. The extracted triples are used to initialize the nodes and determine the edges based on the main relationships between the nodes, forming the final CKG $G$ through the following two steps.
(a) Knowledge graph.
<subject>Frodo
  <predicate>hobbit, own Ring
 <object>Gandalf
  <predicate>guided by
 <object>Sauron
  <predicate>fear, resist
<subject>Gandalf
  <predicate>wizard
 <object>Frodo
  <predicate>ally of
 <object>Sauron
  <predicate>enemy of
<subject>Sauron
  <predicate>Dark Lord, desire Ring
 <object>Frodo
  <predicate>pursue
 <object>Gandalf
  <predicate>enemy of
(b) Linearized knowledge graph.
Figure 3: (a) Part of a knowledge graph generated from The Lord of the Rings, with three named entities. ‘Frodo/Frodo Baggins’ is a single entity with two names. (b) The same graph is in linearized form.
First, to maintain consistency in character identification, we construct a names graph $G_{name}$, consolidating aliases or variations in names in scenes. Our framework processes each scene in turn, extracts all character names, and determines several names refer to same character based on the context using LLM. For example, in The Lord of the Rings, ‘Frodo’ and ‘Frodo Baggins’ are recognized as the same character. As illustrated by ‘Frodo / Frodo Baggins’ in Figure 3(a), each name variation is a node. [^2] In practice, an undirected edges is added between nodes that refer to the same character. This step ensures an accurate capture of relationships, even when names vary across scenes. The CKG is initialized using names from the names graph.
Second, to preserve the consistency of relationships, we sample extracted triples multiple times Wang et al. (2023); Brown et al. (2024) and select frequent ones as the final edges. Let the node set $V$ be the set of all character in the names graph:
$V=\{v\mid v\in G_{\text{name}}\}$ (1)
Only triples with named entities as subjects and objects are used; if an object is missing, a self-loop is added to represent the state of character. We then define the edge set $E$ of our CKG as
$E=\bigl{\{}(s,p,o)\mid s,o\in V,\ \mathrm{freq}(p\mid s,o)\geq\tau\bigr{\}}$ (2)
where $(s,o)$ denotes a character pair, $\mathrm{freq}(p\mid s,o)$ is the frequency of predicate $p$ for $(s,o)$, and $\tau$ is the frequency threshold. [^3] Adjusting the threshold allows for control over the graph: a higher threshold ensures greater consistency, while a lower threshold increases diversity. Finally, the consistent CKG is given by
$G=(V,E).$ (3)
For triples with the same subject and object, frequent predicates capture temporal changes as directed edges. For example, since the early scenes show that ‘Frodo’ fears ‘Sauron’ and the later scenes show that he resists ‘Sauron’, the CKG in Figure 3(a) displays two distinct relationships. The process of deciding edges is repeated to construct a CKG that can effectively evaluate the factuality of summaries.
3.2 NarrativeFactScore Calculation
We invent a new metric to guide agentic evaluation, unlike existing factuality metrics Min et al. (2023); Bishop et al. (2024) that do not provide evidence or feedback for their scores, by considering events in the input story superficially but overlooking relational information about characters. Our metric addresses these limitations by incorporating character relationship graphs and providing detailed feedback. To calculate the factuality of the narrative summary, we first generate an initial summary $Z$ using the prompt described in Section D.2.
To evaluate the factuality of the initial summary $Z$, we decompose it into smaller verifiable units, similar to the approach used in Min et al. (2023). Using the prompt in Section D.3, each sentence in the initial summary $Z$ is divided into a list of atomic facts $A=\{a_{1},a_{2},\ldots,a_{z}\}$.
To evaluate each atomic fact $a_{k}$, we need the scene and information about the characters that appear in the atomic fact. First, we retrieve the most relevant scene $\mathcal{S}_{i}$ within the narrative $\mathcal{N}$, by using the BGE-M3 Chen et al. (2024). Second, we also retrieve the subgraph $g$ from the linearized CKG $G$, as illustrated in Figure 3(b), which contains triples involving the characters mentioned in $a_{k}$.
Using the retrieved information, each atomic fact $a_{k}$ is evaluated to determine its factuality and to obtain feedback supporting the evaluation. We then define an indicator $I_{k}$ for factuality based on:
$I_{k}=\mathbbm{1}[a_{k}\ \text{is factual given}\ \mathcal{S}_{i},g)]$ (4)
where $\mathbbm{1}$ is the indicator function, yielding 1 if the atomic fact $a_{i}$ is factual and 0 otherwise. This evaluation is carried out using the prompt detailed in Section D.4, which produces 1 if the atomic fact is accurate. If the atomic fact is determined to be inaccurate, then feedback $f_{k}$ is provided on how to correct it. Finally, the NarrativeFactScore is calculated as the proportion of atomic facts that are found to be factual, defined by the following equation:
$\text{{NarrativeFactScore}}=\frac{\sum_{i=1}^{z}I_{k}}{z}$ (5)
3.3 Agent-based Fact Refinement
The new metric leveraging consistent CKG enables LLM agent to guide refinement by using feedback from the evaluation. This process involves three key inputs: original narrative to provide global context, the initial summary that requires modification, and the feedback detailing the inaccuracies and reasons for those errors. Using a prompt that incorporates these inputs in Section D.5, LLM refines the initial summary by correcting only the factually inaccurate parts provided from feedback, then generates the improved summary $y$.
Motivated by Madaan et al. (2024), the improved summary can be further evaluated as outlined in Section 3.2. This allows the agent-based refinement to be iterative, where each iteration further refines the summary to enhance overall factuality.
4 Experiments
4.1 Implementation Details
Uniform Language Model Usage
To ensure that performance gain is due to our framework and not the underlying language model, we use only gpt-4o-mini-2024-07-18 in our experiments. This model is applied across all components, including CKG extraction, summarization, fact decomposition, fact check, and agent-based fact refinement. This approach prevents superior LLMs from influencing the results, allowing us to rigorously evaluate the effectiveness of our framework.
Generating Initial Summary
To generate the initial summary $Z$, we adopt hierarchical merging Chang et al. (2024) that ensures the logical structure of the narrative is preserved. The narrative is first divided into chunks $C_{i}$ where each chunk is formed incrementally by adding scenes until a predefined context size444we set predefined context size of a chunk to 1024. is reached. Once this limit is exceeded, a new chunk begins, resulting in a sequence of chunks $\mathcal{C}=\{C_{1},C_{2},...,C_{n}\}$. Each chunk $C_{i}$ is then independently summarized using the prompt specified in Section D.2, and the resulting chunk summaries are sequentially merged to produce the initial summary $Z$.
Retrieving Relevant Scene and Subgraph
Using the BGE-M3 embedding model Chen et al. (2024), we retrieve information relevant to each atomic fact $a_{k}$. Specifically, we identify the most similar scene $\mathcal{S}_{i}$ from the narrative $\mathcal{N}$ and a subgraph containing the three most relevant triples in the linearized CKG $G$. All retrieval computations were performed on a single NVIDIA RTX 3090 GPU.
4.2 Evaluation Metrics
We assess the performance of our framework using several key evaluation metrics. ROUGE Lin (2004) assesses n-gram overlap with reference summaries, including R-1 (unigram), R-2 (bigram) and R-L (longest common subsequence). BERTScore Zhang et al. (2019) (BSp, BSr, BSf1) evaluates similarity using BERT embeddings Devlin et al. (2019), where BSp represents precision, BSr recall and BSf1 the F1-score. BARTScore Yuan et al. (2021) measures the quality of summaries by scoring them as conditional language generation tasks. Finally, we propose NarrativeFactScore (NFS) as a novel metric to measure the factuality of the generated summaries.
We report ROUGE and BERTScore as reference points for lexical and semantic similarity, while emphasizing that these metrics were not designed to capture factual accuracy. Our primary factuality comparisons are instead carried out with dedicated metrics such as FActScore Min et al. (2023) and LongDocFACTScore Bishop et al. (2024).
4.3 Correlation with Human Factuality Scores
Metrics STORYSUMM FABLES
Spearman KENDALL Spearman KENDALL
ROUGE-1 0.25 0.18 -0.20 -0.14
ROUGE-2 0.30 0.22 -0.04 -0.03
ROUGE-L 0.31 0.22 -0.18 -0.14
BERTScoref1 0.19 0.13 -0.13 -0.08
BARTScore 0.09 0.06 -0.30 -0.22
LongDocFACTScore 0.07 0.05 0.24 0.16
FActScore 0.19 0.13 0.16 0.09
NFS 0.43 0.31 0.47 0.33
Table 1: Spearman and KENDALL’s tau correlation coefficients between different metrics and human factuality assessments on STORYSUMM and FABLES. Coefficients indicating strong correlation are underlined. [^6] We follow widely adopted interpretations reported in Table 4.
Metrics STORYSUMM FABLES
Spearman KENDALL Spearman KENDALL
(A) NFS 0.43 0.31 0.47 0.33
(B) 0- consistency 0.21 0.14 0.19 0.13
(C) 0- CKG 0.30 0.21 0.25 0.16
Table 2: Ablation results on STORYSUMM and FABLES, showing the impact of using different CKG.
Dataset
To evaluate whether the NarrativeFactScore we proposed correlates effectively with human factuality, we conducted a series of experiments. For this purpose, we used STORYSUMM  Subbiah et al. (2024) and FABLES Kim et al. (2024), which include multiple summaries generated by LLM for each narrative. These summaries were then evaluated by human annotators based on their factual accuracy.
Results
We computed the Spearman Spearman (1961) correlations and KENDALL’s tau KENDALL (1938) correlations for each metric in relation to the human factuality scores, as shown in Table 6. NarrativeFactScore is the only metric that shows a strong correlation with human annotations in all datasets. This correlation is statistically significant, with p-values below 0.05 for all datasets.
Ablation Study
To verify the effectiveness of CKG in evaluating factuality, we conducted an ablation study in Table 2. NarrativeFactScore (A) iteratively reasons about character relationships, selects frequent relationships to construct a consistent CKG, and utilizes it for factuality evaluation. In contrast, (B) generates the CKG by reasoning character relationships in a single step and evaluates factuality accordingly. However, according to Zhao et al. (2024), LLMs tend to generate inaccurate character relationships when reasoning over long narrative in a single step. Lastly, (C) evaluates factuality without utilizing a CKG.
The experimental results show that our metric (A) achieves the highest correlation with human, and indicate the following observations. First, comparing (A) with (C) shows CKG contributes to more accurate factuality evaluation. However, the results of (B) and (C) show that an inaccurate CKG can hinder factuality evaluation rather than improve it. Thus, to effectively assess the factuality of summary, it is necessary to construct a consistent CKG through multiple iterations of reasoning.
4.4 Summarization Performance Evaluation
MENSA MovieSum
R-1 R-2 R-L BSp BSr BSf1 NFS R-1 R-2 R-L BSp BSr BSf1 NFS
without merging
     TextRank 34.37 4.60 12.84 46.86 49.43 48.10 59.72 33.92 4.62 16.25 46.82 49.48 48.10 60.23
     LED 17.46 1.59 10.03 42.90 42.74 42.58 56.48 2.80 0.28 0.28 32.64 23.82 27.32 22.24
     LongT5 20.77 2.26 10.03 45.05 45.06 45.01 73.76 20.18 1.99 13.83 44.58 44.28 44.36 74.01
hierarchically merging
     GPT-4o-mini 31.79 9.69 12.68 60.00 60.03 60.01 81.05 29.26 8.72 17.88 59.11 59.29 59.19 80.56
     Ours: 1st iteration 33.00 9.70 12.84 60.22 60.11 60.16 85.94 30.36 8.74 18.55 59.26 59.30 59.27 86.92
     Ours: 2nd iteration 33.75 9.72 13.07 60.17 60.10 60.12 88.94 30.98 8.75 18.61 59.33 59.30 59.30 92.04
     Ours: 3rd iteration 34.26 9.74 13.46 60.24 60.21 60.22 93.31 31.31 8.81 18.62 59.36 59.31 59.32 94.59
Table 3: Evaluation results on MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets.
Datasets
We evaluated our framework on the MENSA Saxena and Keller (2024b) and MovieSum Saxena and Keller (2024a) datasets. MENSA aligns movie scenes with Wikipedia summaries and MovieSum pairs screenplays with summaries. We use the full test sets: 50 samples from MENSA and 200 from MovieSum.
Results
We evaluated summarization performance using two baseline types. The first type includes methods without merging that summarize all input in a single step, such as TextRank Mihalcea and Tarau (2004), Longformer Encoder-Decoder (LED) Beltagy et al. (2020), and LongT5 Guo et al. (2022). The second type involves hierarchical merging Chang et al. (2024), with which we performed experiments using GPT-4o-mini OpenAI (2023). Additionally, we evaluated the summaries generated by GPT-4o-mini after agent-based iterative refinements (1st to 3rd).
As shown in Table 3, agent-based refinement improves not only factuality but also other metrics, improving the overall quality of the summaries. This refinement improves performance consistently, yielding improvements of +14.03 in factuality, +2.05 in ROUGE, and +0.13 in BERTScore on MovieSum, and +12.26, +2.47, and +0.21 respectively on MENSA.
5 Analysis
5.1 Analysis for Baseline Metrics
Strength Spearman ($\rho$) KENDALL ($\tau$)
Very weak correlation 0.00~0.15 0.00~0.10
Weak correlation 0.15~0.30 0.10~0.20
Moderate correlation 0.30~0.43 0.20~0.30
Strong correlation 0.43~1.00 0.30~1.00
Table 4: Correlation strength based on coefficients.
[…]
We analyze the results compared to other metrics in Table 6. Metrics based on lexical overlap, such as ROUGE, show stronger correlations with human factuality assessments compared to semantic similarity metrics such as BERTScore, as they better capture repeated entities and locations in narratives. In contrast, metrics such as BARTScore and LongDocFACTScore Bishop et al. (2024), which rely on log-likelihood and entailment, have lower correlations due to their limited ability to account for broader context and character relationships. FActScore Min et al. (2023), reproduced in our study, incorporates character relationship retrieval to improve factuality assessments. Building on this, NarrativeFactScore further enhances performance by addressing common errors caused by misinterpreted character relationships, leading to more accurate evaluations.
5.2 How Consistently Does Ours Capture Character Relationships?
To effectively evaluate factuality and improve summary, it is necessary to generate an accurate and consistent CKG. According to Kim et al. (2024); Zhao et al. (2024), the “naive extract” approach, where an LLM extracts character relationships in one step, often fails to consistently capture some relationships. Thus, our objective is to verify whether our approach can generate a consistent CKG. Conan Zhao et al. (2024) provides ground truth annotation of character relationships within narratives. To evaluate whether the generated relation is semantically similar to this ground truth, we measure the BERTScore Devlin et al. (2019).
As shown in Table 5, our method generates CKG that are closely similar to ground truth. Although the “naive extract” achieves 86.26, it occasionally produces incorrect relationships. In contrast, by reasoning about relationships scene by scene and aggregating them, our method chooses more accurate relationships and constructs a consistent CKG.
Method BSp BSr BSf1
Naive extract Zhao et al. (2024) 86.23 86.33 86.26
Ours 95.63 95.68 95.65
Table 5: Comparison between the naive extract method and our proposed method.
5.3 Challenging Set
R-1 R-2 R-L BSp BSr BSf1 NFS
without merging
     TextRank 33.92 4.63 16.25 46.82 49.48 48.10 62.43
     LED 2.75 0.17 0.64 31.78 24.44 27.37 11.70
     LongT5 22.10 2.29 11.16 43.86 44.69 44.18 79.38
hierarchically merging
     GPT-4o-mini 28.07 8.01 14.12 58.37 59.36 58.53 81.30
     Ours: 1st iteration 29.02 8.09 14.08 58.49 59.30 58.86 84.59
     Ours: 2nd iteration 29.98 8.19 14.24 58.61 59.32 58.94 90.47
     Ours: 3rd iteration 30.22 8.20 14.44 58.75 59.39 59.04 93.22
Table 6: Evaluation results on the challenging set of the MovieSum Saxena and Keller (2024a) dataset.
We aim to evaluate whether our metric can provide feedback necessary to improve factuality in recent narratives. Although LLM-based metrics provide accurate factuality feedback for narratives within their pretraining data, they often fail for narratives outside of their training corpus. However, our metric provides accurate feedback by evaluating summaries based on narrative story and character relationships rather than relying on parametric knowledge alone. Therefore, we define a challenging set of works published after the knowledge cutoff date of our LLM to verify whether our metric improves factuality through its feedback.
Our metric demonstrates the capability to provide feedback for improving factuality even in recent works. For this experiment, we curated a challenging set of 18 movies from MovieSum Saxena and Keller (2024a) released after our LLM knowledge cutoff. [^8] We used GPT-4o-mini with an October 2023 knowledge cutoff date. We conducted refinement experiments identical to Table 3 to correct factual errors in this challenging set. As shown in Table 6, three rounds of refinement improved NarrativeFactScore by 11.92, comparable to the improvements in Table 3. These results confirm our metric provides effective feedback for recent stories independent of LLM parametric knowledge.
5.4 Latency
Metric
CKG
extraction
time
Factuality
calculation
time
KENDALL
Human - 132.00 min -
LongDocFACTScore - 3.81 min 0.16
FActScore - 4.60 min 0.09
NarrativeFactScore 1.17 min 4.81 min 0.33
Table 7: Average latency (in minutes) and KENDALL’s tau correlation for evaluating factuality across different metrics on the FABLES Kim et al. (2024) dataset.
We compared the latency of LLM-as-a-Judge metrics, such as LongDocFACTScore and FActScore, with our Agent-as-a-Judge metric. Table 7 shows the time required to evaluate the factuality of each summary in the FABLES Kim et al. (2024). Since long narratives like those in FABLES exceed 100K tokens, human evaluation by verifying details is time-consuming. [^9] According to Kim et al. (2024), annotators spent over 11 hours evaluating five summaries. In contrast, LLM-based metrics, including ours, assess factuality within a few minutes. However, LLM-as-a-Judge metrics struggle to assess factuality while understanding character relationships, leading to discrepancies with human evaluations, as shown in Table 6. In contrast, NarrativeFactScore devotes additional time to reasoning about character relationships before assessing factuality, resulting in more accurate evaluations despite slightly longer times.
5.5 Sensitivity of Metrics on Factual Perturbation
Summary ROUGE-L BERTScoref1 NFS
Reference Summary 100.00 100.00 95.42
Perturbed Summary 81.61 92.15 40.81
Table 8: Change in metric scores after factual perturbation of the reference summary on the MENSA.
We evaluated metric sensitivity to factual perturbations using GPT-4o with the prompt shown in Figure 10 on the MENSA. Specifically, we perturbed the reference summaries from MENSA by introducing factual inaccuracies in each sentence. Table 8 shows that ROUGE-L and BERTScoref1 decreased minimally despite factual perturbations, while NarrativeFactScore significantly dropped. These results demonstrate that our metric is highly sensitive to factual discrepancies, making it a suitable metric for assessing factuality.
6 Conclusion
This work shows how the agent-as-judge is deployed to overcome the limitations of existing evaluation metrics, such as overreliance on lexical similarity or factual inconsistencies. Specifically, we propose consistent CKG extraction, and new factual evaluation metric based on CKG, and an agent that evaluates and guides the summary and refinement. Through our implementation, we demonstrated both the process and superior performance over state-of-the-art methods on real-life industry datasets and scenarios.
7 Limitation
We acknowledge two limitations of our work. First, our framework may occasionally retrieve subgraphs unrelated to the atomic fact being evaluated, though this did not impact factuality judgments in our experiments and outperformed the no-retrieval baseline. Nonetheless, further enhancing subgraph retrieval precision remains a promising direction for future work.
Second, our framework has been tested exclusively in the narrative domain. Although effective, its generalizability to other domains remains unverified. However, its potential for applications requiring deep character understanding—such as news summarization, biographical writing, and historical analysis—suggests promising directions for future exploration.
Appendices
Appendix A Qualitative Example
Initial
summary
Saruman reveals to Gandalf that Sauron is regaining strength and is gathering an army to attack Middle-earth, using a Palantir to see his plans. Gandalf warns Saruman about the dangers of the Palantir and expresses disbelief at Saruman’s willingness to ally with Sauron, leading to a fierce battle between the two. In the subsequent scene, Frodo and Sam are enjoying a peaceful moment in the Shire when they are interrupted by Merry and Pippin, who crash into them after stealing vegetables from Farmer Maggot’s field, prompting a humorous chase as they flee from the angry farmer and his dogs. The outcome sees Gandalf and Saruman in conflict, while Frodo and his friends are caught up in a lighthearted escapade.
Atomic
facts
[1] Saruman reveals to Gandalf that Sauron is regaining strength.
[2] Sauron is gathering an army to attack Middle-earth.
[3] Sauron uses a Palantir to see his plans.
[4] Gandalf warns Saruman about the dangers of the Palantir.
[5] Gandalf expresses disbelief at Saruman’s willingness to ally with Sauron.
[6] A fierce battle occurs between Gandalf and Saruman.
[7] Frodo and Sam enjoy a peaceful moment in the Shire.
[8] Merry and Pippin crash into Frodo and Sam.
[9] Merry and Pippin steal vegetables from Farmer Maggot’s field.
[10] A humorous chase ensues as they flee from Farmer Maggot and his dogs.
[11] Gandalf and Saruman are in conflict.
[12] Frodo and his friends are caught up in a lighthearted escapade.
Factuality
calculation
[1] True
[2] True
[3] False, The statement is false because Sauron does not use a Palantir to see his plans; rather, it is Saruman who uses the Palantir to gain knowledge about Sauron’s actions and intentions.
(evidence scene: #39, evidence subgraph: ‘Saruman-own-Palantir’, …)
[4] True
[5] True
[6] True
[7] False, The statement “Frodo and Sam enjoy a peaceful moment in the Shire” is false. The scene depicts Frodo and Sam being interrupted by Merry and Pippin, leading to a messy situation as they are chased by Farmer Maggot and his dogs after stealing vegetables from his field.
(evidence scene: #40, evidence subgraph: ‘Frodo-friend-Sam’, …)
[8] True
[9] True
[10] True
[11] True
[12] True
Refined
summary
Saruman reveals to Gandalf that Sauron is regaining strength and is gathering an army to attack Middle-earth. Saruman has used a Palantir to gain insight into Sauron’s plans. Gandalf warns Saruman about the dangers of the Palantir and expresses disbelief at Saruman’s willingness to ally with Sauron, leading to a fierce battle between the two. In the subsequent scene, Frodo and Sam are caught in a messy situation in the Shire when Merry and Pippin crash into them after stealing vegetables from Farmer Maggot’s field, prompting a humorous chase as they flee from the angry farmer and his dogs. The outcome sees Gandalf and Saruman in conflict, while Frodo and his friends are caught up in a lighthearted escapade.
Table 9: Qualitative example illustrating NarrativeFactScore. Red text in the initial summary and atomic facts indicate factually incorrect statements based on scene evidence, while blue text in the refined summary indicate corrections made through agent-based refinement based on feedback.
In this section, we illustrate how our approach evaluates factuality, provides actionable feedback, and refines summaries through a qualitative example. Table 9 shows the evaluation and refinement process for a summary of The Lord of the Rings generated by GPT-4o-mini OpenAI (2023). The CKG has already been constructed using the method described in Section 3.1.
To evaluate factuality, we decompose the initial summary into atomic facts. The summary contains two factually incorrect statements highlighted in red, which are also presented in the atomic facts. First, according to the original script, Saruman uses a Palantir to observe Sauron; however, due to difficulty in understanding character relationships, the incorrect summary stating Sauron observes Saruman was generated and recorded as atomic fact [3]. Second, while the original script depicts Frodo and Sam in a messy situation, chased in the Shire by Merry and Pippin, the summary incorrectly describes it as peaceful, recorded as atomic fact [7].
In our framework, we retrieve relevant scene and subgraph for each atomic fact to evaluate factuality. Consequently, only [3] and [7] were identified as false. For these facts, the framework generates not only the factuality but also actionable feedback explaining why they are false and how to correct them. For [3], based on retrieved scene and relationship that Saruman owns the Palantir, our framework determines that the statement is false and generates feedback suggesting that Saruman uses the Palantir to gain knowledge about Sauron’s actions and intentions. Similarly, for [7], based on the scene showing Frodo and Sam in a messy situation with Merry and Pippin in the Shire, our framework determines the statement is false and provides proper feedback. Using this detailed feedback, our framework generates refined summary by correcting only the erroneous parts of the initial summary.
Appendix B System Deployment
Figure 4: Deployment overview of NarrativeFactScore.
This section describes our system deployment, which is necessary for the media industry [^10] Company name anonymized for blind reviewing. The media industry broadly refers to the sector that creates, distributes, and analyzes various forms of narratives, such as movies, television shows, books, video games, and other media that tell stories. This includes businesses involved in producing, editing, and consuming these forms of content, focusing on storytelling in both traditional and digital media. where companies make investment decisions on narratives (> 100K tokens) such as dramas or movies. Since reading all narratives is challenging, media companies utilize summaries of each narrative to determine its production feasibility. However, summaries generated by humans or LLMs frequently contain factual inconsistencies, which hinder accurate investment decisions. Therefore, our proposed system is deployed to evaluate the factuality of summary for long narratives and improve its factuality.
Figure 4 shows screenshots of the system111111bit.ly/4iz8pXc, aligned with the three phases of our framework in Figure 2. Using the example Black Panther, users can view the original narrative after selecting a dataset, data type, and name. Clicking “Generate Knowledge Graph” generates and visualizes the CKG (Section 3.1). The “Generate Initial Summary” and “Calculate Factuality Score” buttons create an initial summary and evaluate its factuality using the CKG (Section 3.2). Finally, “Refine Summary” improves the summary based on feedback, enhancing factuality (Section 3.3).
Appendix C Usage of AI Assistants
We utilized ChatGPT to improve the clarity and grammatical accuracy of my writing. It provided suggestions for rephrasing sentences and correcting grammatical errors to make the text flow more naturally.
Appendix D Prompts
To ensure ethical transparency and reproducibility, we disclose the prompts used at each stage of our process.
D.1 Knowledge Extraction Prompt for LLM
Knowledge Extraction Prompt [Begin story excerpt] “Christmas won’t be Christmas without any presents,” grumbled Jo. “It’s so dreadful to be poor!” sighed Meg, looking out the window at the snow-covered streets of Concord. “I don’t think it’s fair…” “Glad to find you so merry, my girls,” said a cheery voice at the door… “A letter! A letter! Three cheers for Father!” [End story excerpt] Named entities: Jo / Jo March Meg / Margaret / Margaret March Amy Beth / Elizabeth March sisters Mrs. March / Marmee / Mother Father Concord Union Army Knowledge graph edges: 1. Jo, Meg, Amy, Beth; in; March sisters 2. March sisters; daughters of; Mrs. March, Father 3. Mrs. March; mother of; March sisters 15. Mrs. March; brought home a letter from; Father [Begin story excerpt] {scene of narrative} [End story excerpt]
Figure 5: Simplified prompt for named entity recognition and knowledge graph edges generation.
D.2 Narrative Summarization Prompt for LLM
Narrative Summarization Prompt This is a part of a script from a Movie. Read the following content carefully, then answer my question: {chunk of narrative} The script has ended now. Summary instructions: - Provide a detailed summary of the key characters’ actions, emotions, and situations as reflected in the dialogue or context. - Clearly state the outcome of the events. - The summary should be between 2 to 5 sentences long.
Figure 6: Prompt for summarizing a chunk of narrative from a movie script.
D.3 Atomic Fact Decomposition Prompt for LLM
[…]
Figure 7: Prompt for extracting atomic facts from a movie script summary.
D.4 Fact-Checking Prompt for NarrativeFactScore
Fact-Checking Prompt Consider the given statement, the related scene, and the relationship subgraph. Indicate whether the statement is supported by the scene and the relationship subgraph. Negation of a false statement should be considered supported. If the statement is true, output 1. If the statement is false, output the reason why it is false. Scene: {retrieved scene} Relationship Subgraph: {retrieved subgraph} Statement: {atomic fact} Output:
Figure 8: Prompt for validating a summary against a scene and a relationship subgraph.
D.5 Agent-based Refinement Prompt for LLM
Agent-based Refinement Prompt Below is a part of the script from the titled movie. - Script: {chunk of narrative} Based on the ’Statement to Revise’ and ’Reason for Revision’, create a ‘Revised Summary’ of the ‘Summary of the Script’. Keep the revised summary concise and similar in length to the original summary. Do not directly copy any part of the ’Script.’ If the ’Summary of the Script’ is accurate, generate the original summary as is. - Summary of the Script: {initial summarization} - Statement to Revise 1: {hallucinated fact atomic} (Reason for Revision: {feedback}) - Revised Summary:
Figure 9: Prompt for revising and summarizing a movie script based on feedback. Note that ‘Statement to Revise’ and ‘Reason for Revision’ correspond to the atomic fact and factuality feedback calculated in Figure 8.
D.6 Factual Perturbation Prompt
Narrative Summarization Prompt This sentence serves as a summary of a script. Rewrite this one-sentence summary by minimally replacing a few words in the original sentence to render it factually inaccurate, while keeping the original sentence structure intact. Original sentence: {original_sentence} Rewritten sentence:
Figure 10: Prompt used to generate factual perturbations.